Please explain with proper steps.

If the variance of the first n natural numbers is 10 and the variance of the first m even natural numbers is 16, then m+n is equal to _____.

Grade:12

2 Answers

Ram Kushwah
3 years ago
Dear student ,
The variant of  natural numbers from 1 to n is given by
(1²+2²+3²+...................+n²)/n - (1+2+3+...........+n)²
 
In the given equation we have fount n=11 and m=7
 
Thus m+n=11+7=18
Thanks
 
 
Ram Kushwah
3 years ago
Please delete my first answer 
 
Dear student ,
The variant of  natural numbers from 1 to n is given by
(1²+2²+3²+...................+n²)/n - (1+2+3+...........+n)²
 
In the solution of the given question we have found n=11 and m=7
 
Thus m+n=11+7=18
Thanks